Mrs. Singer owns a profitable sole proprietorship. For each of the following cases, use a Schedule SE, Form 1040, to compute her 2017 self-employment tax and her income tax deduction for such tax.

  1. Mrs. Singer’s net profit from Schedule C was $51,458. She had no other earned income.

  2. Mrs. Singer’s net profit from Schedule C was $51,458, and she received a $80,000 salary from an employer.

  3. Mrs. Singer’s net profit from Schedule C was $51,458, and she received a $137,000 salary from an employer.
Answer #1

1) In First Case, When Mrs. Singer’s net profit from Schedule C was $51,458. She had no other earned income.

Self- employment tax will be = $ 51458 * 92.35% * 15.3% (Since $ 51458 * 92.35% is less than $ 127200)

= $ 7270.78

Income tax deduction = $ 7270 / 2 = $ 3635.39

2) When Mrs. Singer’s net profit from Schedule C was $51,458, and she received a $80,000 salary from an employer.

Total earnings = $ 51458 + $ 80000 = $ 131458

Self- employment tax will be = $ 131458 * 92.35% * 15.3% (Since $ 131458 * 92.35% is less than $ 127200)

= $ 18574.42

Income tax deduction = $ 18574.42 / 2 = $ 9287.21

3) When Mrs. Singer’s net profit from Schedule C was $51,458, and she received a $137,000 salary from an employer.

Total earnings = $ 51458 + $ 137000 = $ 188458

Self- employment tax will be = $ 188458 * 92.35% * 2.9%% (Since $ 188458 * 92.35% is more than $ 127200) + $ 15772.80

= $ 20820

Income tax deduction = $ 20820 / 2 = $ 10410